# Towards a complete scheme of cosmological neutrino self-interactions: Collision term for a wide range of mediator masses

### v1

@@ 3.3 $\beta_{\ell}$ coefficients for the effective massless neutrino ($m_{i}\approx 0$)
$\frac{\partial{\cal F}_{\ell}}{\partial\tau}+k\left[\frac{(\ell+1)}{(2\ell+1)}{\cal F}_{\ell+1}-\frac{\ell}{(2\ell+1)}{\cal F}_{\ell-1}\right]-4f_{g}=-a\Gamma_{\ell}(\mu_{\varphi}){\cal F}_{\ell}\,.$ (3.49)
@@
$\Gamma_{\ell}(\mu_{\varphi})=\frac{g^{4}T_{\nu}}{16}\begin{cases}{\rm g}_{s}\left[\frac{1}{2}\beta_{\ell}(\mu_{\varphi})+\sum_{j\neq i}\beta_{\ell}^{j}(\mu_{\varphi})\right]&\text{for Majorana-like neutrinos}\\ \frac{1}{2}\beta_{\ell}(\mu_{\varphi})+\sum_{j\neq i}\beta_{\ell}^{j}(\mu_{\varphi})&\text{for Dirac-like (anti)neutrinos}\end{cases}\,.$ (3.50)
@@
$\beta_{\ell}^{j}(\mu_{\varphi}):=\frac{15}{7\pi^{7}}\int_{0}^{\infty}dq_{1}q_{1}^{2}\,{\cal K}_{\ell}^{ij}(q_{1})\,.$ (3.51)
@@
$\beta_{\ell}(\mu_{\varphi}):=\frac{15}{7\pi^{7}}\int_{0}^{\infty}dq_{1}q_{1}^{2}\,{\cal K}_{\ell}(q_{1})\,.$ (3.52)
@@
Notice that the functions (3.51) and (3.52) implicitly depend on the reference temperature, since $\mu_{\varphi}=m_{\varphi}/T_{\nu}$ and $\mu_{j}=m_{j}/T_{\nu}$. Also, given that the monopole and dipole moments do not contribute to the collision term $\beta_{0,1}^{(j)}=0$. This will serve as a test to verify that our numerical calculation is working correctly.
@@ 4.1 The heavy mediator limit
The amplitudes take the form (E.29) and (E.30), with the terms $A$, $B_{\ell}$ and $D_{\ell}$ given by the integrals summarized in Eqs. (E.31)–(E.33). To establish the previous results, we used that $y_{-}=x_{-}=1$ in the limit of zero neutrino mass from (3.24) and (3.23). Using our approach, we have verified the accuracy of the HML at the level of $\mu_{\varphi}\gtrsim 10^{2}$ with desviations $\lesssim 0.6\%$, in this regime, it is a good approximation even for massive neutrinos.
@@ 4.2 Beyond the heavy mediator limit
Figure 5 provides a quantitative insight into the relevance of the resonance. The resonant regime originates from the presence of contributions of the channel $s$; this resonant region, as we preliminarily studied, lies roughly in the region given by $\mu_{\varphi}\sim 2\times 10^{-3}-3\times 10^{1}$, for effectively massless neutrinos. In Fig. 5, recall that we explicitly subtract the resonant contribution from the $\beta_{\ell}^{(j)}$ functions by applying the regularization scheme of Eqs. (3.46), (3.34) and (3.33) in which the $s$-channel contribution is set to zero for $B_{x}^{\prime}=0$ (or $B_{x,ij}^{\prime}=0$).
@@
Modeling the resonant region is difficult; however, it has been approximated in Refs.  [112, 96], and in other contexts for astrophysical neutrinos  [95, 27, 38, 59, 115]. Studying in detail the resonant region is beyond the scope of this manuscript, and we expect to give full details in future work.
@@ 5 Light mediator limit vs relaxation time approximation
For the light mediator, $m_{\varphi}\lesssim 10^{-3}\,{\rm eV}$, studies of $\nu-\nu$ scattering [74] have used the RTA, in which the effect of interactions is modeled by driving the distribution function $f_{\nu_{i}}$ toward its equilibrium form $f_{\nu_{i}}^{(0)}$ over a characteristic timescale $t_{c}$. In this approximation, the collision term takes the simple form:
@@
$\Gamma_{\nu\nu\to\nu\nu}=\frac{{\rm g}_{s}}{2\pi^{2}}\frac{3}{2}\zeta(3)g_{\rm eff}^{4}T_{\nu}\,.$ (5.2)
@@
In Table 1, we show the first $\xi_{\ell}$ values up to $\ell=10$, we observe a saturation for $\ell\geq 20$ and differences smaller than $1\,\%$. Furthermore, with the convention in Dirac-like neutrinos (degenerate in neutrino/antineutrino mass), we have $\xi_{\ell}\,({\rm Majorana})\approx 2\xi_{\ell}\,$(Dirac-like). This factor is due to $s$-channel suppression.
@@
The current constraints on a light mediator for Majorana neutrinos in $\nu-\nu$ scattering require an effective coupling $g_{\rm eff}<1.94\times 10^{-7}$, based on Planck 2018 data [113], which translate for only one process to $g<9.52\times 10^{-7}$ for $\ell=2$ or $g<8\times 10^{-7}$ for $\ell=20$, reflecting that $\xi_{\ell}^{1/4}\sim\mathcal{O}(10^{-1})$. However, these bounds can be updated by including a complete collision term that accounts for additional processes, as $\nu_{i}\nu_{j}\to\nu_{i}\nu_{j}$, as shown in Eq. (3.50); moreover, we showed that the contributions are $\ell$-dependent.
@@ 6 Summary and conclusions
We started reviewing the Boltzmann hierarchy equations for massive neutrinos. In the following sections, we computed our most general form of the collision term, which in general distinguishes between mass eigenstates and neutrinos being Dirac-like or Majorana particles. Furthermore, our framework includes both the neutrino mass and the mediator mass as free parameters. From there, we systematically distill the three approximations used commonly in the literature: the light, heavy, and resonant mediators. In each case, we compared our results with previous results and showed the advantages and limitations of our approach.
@@
The commonly used HML is a popular framework that has the potential to address cosmological tensions. However, this approach has previously relied on the effective coupling $G_{\rm eff}$ combined with a tight coupling approximation at high temperatures. The latter is particularly well motivated, as previous computations did not permit a systematic exploration of the coupling strength and the mediator mass. In contrast, our work allows exploration of the parameter space using the fundamental coupling $g$ directly, enabling explicit tests of the mediator mass $m_{\varphi}$.
@@
Here, we showed that the neutrino mass does not affect the Boltzmann collision term for $m_{\nu}\lesssim 0.06\,{\rm eV}$ for redshifts larger than $1.8\times 10^{4}$. We also relaxed the heavy mediator validity condition down to $m_{\varphi}\geq 10^{2}\,{\rm eV}$, and showed that Dirac-like neutrinos do not exhibit resonance production in $\nu-\nu$ scattering, but they do in the $\nu_{i}-\bar{\nu}_{j}$ scattering. Modeling the resonant regime for Majorana neutrinos remains challenging since it requires the inclusion of one-loop corrections for a complete amplitude description. We anticipate that a full understanding of these corrections will allow for a thorough exploration of heavy mediators in NSI in future work. Finally, we provide an exact treatment for the light mediator regime, establishing that for $m_{\varphi}\lesssim 10^{-3}\,{\rm eV}$ the mediator mass and neutrino mass ($m_{j}\lesssim 0.06\,{\rm eV}$) can be safely neglected according to our numerical analysis for $z\gtrsim 6\times 10^{3}$ ($T_{\nu}\gtrsim 1\rm{eV}$). To conclude, we were able to find the correct value of $\xi_{l}$, which is typically codified into $g_{\rm eff}=\xi^{1/4}g$ that was not previously computed, and find the multipolar dependence on $\ell$. All of our results would be useful for future cosmological data analysis of neutrino self-interactions.
@@
Another important feature is that interactions such as (3.1) can produce scalar particles through the process $\varphi$, via $\nu\nu\to\varphi\varphi$ and $\nu\nu\to\varphi$ when the temperature is comparable to $m_{\varphi}$. This may invalidate the assumption $C^{(0)}=0$, or at least render its contribution subdominant rather than negligible. We plan to address this effect and provide a more accurate description in future work.
@@ Acknowledgments
The metric perturbation $h_{ij}$ can be decomposed into a trace part and a traceless part; The trace part is defined as $h={h^{i}}_{i}$, and the traceless part consists of three components: $h^{(k)}_{ij}$, $h^{(\perp)}_{ij}$, and $h^{(T)}_{ij}$. Thus, we write $h_{ij}=h\,\delta_{ij}/3+h^{(k)}_{ij}+h^{(\perp)}_{ij}+h^{(T)}_{ij}$. By definition, the divergences of $h^{(k)}_{ij}$ and $h^{(\perp)}_{ij}$ correspond to longitudinal and transverse vectors, respectively, while $h^{(T)}_{ij}$ is fully transverse. The components of $h_{ij}$ in Eq. (2.1) satisfy the following conditions: $\epsilon_{ijk}\,\partial_{j}\partial^{l}h^{(k)}_{lk}=0$, $\partial^{i}\partial^{j}h^{(\perp)}_{ij}=0\,,\partial^{i}h^{(T)}_{ij}=0\,.$ It follows that $h^{(k)}_{ij}$ can be expressed in terms of a scalar field $\chi$, and $h^{(\perp)}_{ij}$ in terms of a divergenceless vector field $A_{i}$, as
$h^{(k)}_{ij}=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\chi\,,h^{(\perp)}_{ij}=\partial_{i}A_{j}+\partial_{j}A_{i}\,,\quad\partial^{i}A_{i}=0\,.$ (A.1)
@@ Appendix B Collision integral reduction for $\alpha-\beta$ elastic scattering processes
$F^{(1)}=\Psi_{S}\left[2K_{3}\Theta({\bf p}_{3})-K_{1}\Theta({\bf p}_{1})-K_{2}\Theta({\bf p}_{2})\right]$ (B.2)
@@
for different particles in the final state. Where we have dropped an implicit $\tau$ and the $\bf x$} dependence, here the $K_{n}$-terms are determined by
@@
The general form of $\Psi_{S}$ is valid for any elastic scattering processes $\alpha(P_{1})+\beta(P_{2})\rightarrow\alpha(P_{4})+\beta(P_{3})$ and annihilation processes as $\alpha+\alpha\to\beta+\beta$, since the necessary and sufficient condition to establish this expression is $\sigma_{1}\sigma_{2}=\sigma_{3}\sigma_{4}$, though, we will not analyze ineslastic processes here  [^15] In certain inelastic scattering processes, the same condition on sigmas applies, such as in $\nu_{i}\nu_{j}\to\varphi\varphi$ or $\nu_{i}\nu_{j}\to\nu_{k}\nu_{l}$.. Note that the double contribution of $\Theta({\bf p}_{3})$ in Eq.(B.2) comes from $\Theta({\bf p}_{4})$ after the explicit use of the ${\bf p}_{4}\leftrightarrow{\bf p}_{3}$ symmetry for identical particles in the final state. The corresponding rescaled collision term in Fourier space then behaves as follows
@@
$\frac{d^{3}p_{4}}{2E_{4}}=d^{4}P_{4}\delta(P_{4}^{2}-m_{j}^{2})H(E_{4}-m_{j})\quad\text{and}\quad\frac{d^{3}p_{3}}{2E_{3}}=d^{4}P_{3}\delta(P_{3}^{2}-m_{i}^{2})H(E_{3}-m_{i})\,.$ (B.11)
@@
$(P_{1}+P_{2}-P_{4})^{2}-m_{i}^{2}$ $=2\left(P_{1}\cdot P_{2}-P_{1}\cdot P_{4}-P_{2}\cdot P_{4}+m_{j}^{2}\right)\,.$
because of Eq. (B.12), fixes the angle to the value $\beta_{0}$ ($\beta_{0}^{\prime}$) that comply with the condition
@@
$\cos\beta_{0}^{\prime}$ $=\frac{E_{2}E_{4}-E_{1}(E_{2}-E_{4})-m_{j}^{2}+p_{1}p_{2}\cos\alpha-p_{1}p_{4}\cos\theta^{\prime}-p_{2}p_{4}\cos\alpha\cos\theta^{\prime}}{p_{2}p_{4}\sin\alpha\sin\theta^{\prime}}\,.$ (B.15)
@@
$\sin^{2}\beta_{0}=\frac{\tilde{S}(x,y)}{|p_{2}p_{3}\sin\alpha\sin\theta|^{2}}\quad\text{and}\quad\sin^{2}\beta_{0}^{\prime}=\frac{\tilde{S}^{\prime}(x,z)}{|p_{2}p_{4}\sin\alpha\sin\theta^{\prime}|^{2}}\,.$ (B.19)
@@
    $c_{x,j}^{\prime}$ $=-[z_{2}z_{4}+z_{1}(z_{4}-z_{2})-\mu_{j}^{2}+q_{1}q_{2}x]^{2}+q_{2}^{2}q_{4}^{2}(1-x^{2})~,$ (B.28)
    or
  • $S^{\prime}(x,z)=a_{z}x^{2}+b_{z,j}x+c_{z,j}$, where
@@
with $I_{n}$ are the angular integrals
$I_{n}=\begin{cases}I_{n}(\mu_{n},p_{1},p_{2},p_{3},\tau)=\frac{1}{T^{2}}\int\frac{dxdy\,d\phi}{\sqrt{S(x,y)}}\Theta_{n}({\bf k}.{\bf p}_{n},\tau)\overline{|{\cal M}|}^{2}H[S(x,y)]\quad\text{for $n=1,2,3$}\\ I_{4}(\mu_{4},p_{1},p_{2},p_{4},\tau)=\frac{1}{T^{2}}\int\frac{dxdz\,d\phi^{\prime}}{\sqrt{S^{\prime}(x,z)}}\Theta_{4}({\bf k}.{\bf p}_{4},\tau)\overline{|{\cal M}|}^{2}H[S^{\prime}(x,z)]\quad\text{for $n=4$}\end{cases}\,.$ (B.33)
Both were evaluated in the momentum conservation condition $P_{4}=P_{1}+P_{2}-P_{3}$. After inserting the Legendre expansion (2.14), where $k=|{\bf k}|$, and $P_{\ell}$ is the Legendre polynomial of order $\ell$,
@@
${\cal J}_{4}(p_{1},\tau)$ $=\frac{1}{8(2\pi)^{4}E_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1){\cal P}_{\ell}(\upmu)\,\int\!\frac{p_{2}^{2}dp_{2}}{E_{2}}\frac{p_{4}^{2}dp_{4}}{E_{4}}\Psi_{S}K_{4}H(E_{3}-m_{i})\vartheta_{l}(p_{4})\tilde{\cal I}_{4l}\,.$
@@
${\cal I}_{4\ell}(q_{1},q_{2},q_{4})$ $=\int\frac{dxdz}{\sqrt{S^{\prime}}}\overline{|{\cal M}|}^{2}\,H(S^{\prime})\,{\cal P}_{\ell}(z)\,.$
@@
$\tilde{C}^{(1)}[{\bf k},p_{1},\tau]=\frac{1}{4\pi}\frac{T_{\nu}}{128\pi^{3}z_{1}}\frac{d\ln f^{(0)}(p_{1})}{d\ln p_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1)\nu_{\ell}{\cal P}_{\ell}(\upmu)\tilde{{\cal K}}(z_{1})\,.$ (B.38)

### v2

@@ 3.3 $\beta_{\ell}$ coefficients for the effective massless neutrino ($m_{i}\approx 0$)
$\frac{\partial{\cal F}_{\ell}}{\partial\tau}+k\left[\frac{(\ell+1)}{(2\ell+1)}{\cal F}_{\ell+1}-\frac{\ell}{(2\ell+1)}{\cal F}_{\ell-1}\right]-4f_{g,\ell}=-a\Gamma_{\ell}(\mu_{\varphi}){\cal F}_{\ell}\,.$ (3.49)
@@
$\Gamma_{\ell}(\mu_{\varphi})=\frac{g^{4}T_{\nu}}{16}\begin{cases}{\rm g}_{s}\left[\frac{1}{2}\beta_{\ell}(\mu_{\varphi})+\sum_{j\neq i}\beta_{\ell}^{j}(\mu_{\varphi})\right]&\text{for Majorana-like neutrinos}\\ \frac{1}{2}\beta_{\ell}(\mu_{\varphi})+\sum_{j\neq i}\beta_{\ell}^{j}(\mu_{\varphi})&\text{for Dirac-like (anti)neutrinos}\end{cases}\,$, (3.50)
@@
$\beta_{\ell}^{j}(\mu_{\varphi}):=\frac{15}{7\pi^{7}}\int_{0}^{\infty}dq_{1}q_{1}^{2}\,{\cal K}_{\ell}^{ij}(q_{1})\,$, (3.51)
@@
$\beta_{\ell}(\mu_{\varphi}):=\frac{15}{7\pi^{7}}\int_{0}^{\infty}dq_{1}q_{1}^{2}\,{\cal K}_{\ell}(q_{1})\,$, (3.52)
@@
Notice that the functions (3.51) and (3.52) implicitly depend on the reference temperature, since $\mu_{\varphi}=m_{\varphi}/T_{\nu}$ and $\mu_{j}=m_{j}/T_{\nu}$. Also, given that the monopole and dipole moments do not contribute to the collision term $\beta_{0,1}^{(j)}=0$. This will serve as a benchmark to verify that our numerical implementation using VEGAS algorithm [105, 88] is working correctly.
@@ 4.1 The heavy mediator limit
The amplitudes take the form (E.27) and (E.28), with the terms AA, $B_{\ell}$ and $D_{\ell}$ given by the integrals summarized in Eqs. (E.29)–(E.31). To establish the previous results, we used that $y_{-}=x_{-}=1$ in the limit of zero neutrino mass from (3.24) and (3.23). Using our approach, we have verified the accuracy of the HML at the level of $\mu_{\varphi}\gtrsim 10^{2}$ with desviations $\lesssim 0.6\%$, in this regime, it is a good approximation even for massive neutrinos.
@@ 4.2 Beyond the heavy mediator limit
Figure 5 provides a quantitative insight into the relevance of the resonance. The resonant regime originates from the presence of contributions of the channel $s$; this resonant region, as we preliminarily scanned, lies roughly in the region given by $\mu_{\varphi}\sim 2\times 10^{-3}-5\times 10^{1}$, for effectively massless neutrinos. In Fig. 5, recall that we explicitly subtract the resonant contribution from the $\beta_{\ell}^{(j)}$ functions by applying the regularization scheme of Eqs. (3.46), (3.34) and (3.33) in which the $s$-channel contribution is set to zero for $B_{x}^{\prime}=0$ (or $B_{x,ij}^{\prime}=0$).
@@
Modeling the resonant region is difficult; however, it has been approximated in Refs.  [116, 99], and in other contexts for astrophysical neutrinos  [98, 28, 39, 60], and has recently gained relevance in scenarios where the lightest neutrino species remains relativistic today [119, 92]. Studying in detail the resonant region is beyond the scope of this manuscript, and we expect to give full details in future work.
@@ 5 Light mediator limit vs relaxation time approximation
For the light mediator, $m_{\varphi}\lesssim 10^{-3}\,{\rm eV}$, studies of $\nu-\nu$ scattering [75] have used the RTA, where the effect of interactions is modeled by driving the distribution function $f_{\nu_{i}}$ toward its equilibrium form $f_{\nu_{i}}^{(0)}$ over a characteristic timescale $t_{c}$. In this approximation, the collision term takes the simple form:
@@
$\Gamma_{\nu\nu\to\nu\nu}=\frac{{\rm g}_{s}}{2\pi^{2}}\frac{3}{2}\zeta(3)g_{\rm eff}^{4}T_{\nu}\,$, (5.2)
@@
In Table 1, we show the first $\xi_{\ell}$ values up to $\ell=10$, we observe a saturation for $\ell\geq 20$ and differences are smaller than $1\,\%$. Furthermore, with the convention in Dirac-like neutrinos (degenerate in neutrino/antineutrino mass), we have $\xi_{\ell}\,({\rm Majorana})\approx 2\xi_{\ell}\,$(Dirac-like). This factor is due to $s$-channel suppression.
@@
The current constraints on a light mediator for Majorana neutrinos in $\nu-\nu$ scattering require an effective coupling $g_{\rm eff}<1.94\times 10^{-7}$, based on Planck 2018 data [117], which translate for a single process to $g<9.52\times 10^{-7}$ for $\ell=2$ or $g<8\times 10^{-7}$ for $\ell=20$, reflecting that $\xi_{\ell}^{1/4}\sim\mathcal{O}(10^{-1})$. However, these bounds can be updated by including a complete collision term that accounts for additional processes, as $\nu_{i}\nu_{j}\to\nu_{i}\nu_{j}$, as shown in Eq. (3.50); moreover, we showed that the contributions are $\ell$-dependent.
@@ 6 Summary and conclusions
Along the present paper, we have revisited Boltzmann hierarchy equations for massive neutrinos and computed our most general form of the collision term, which in general distinguishes between mass eigenstates and whether neutrinos are Dirac-like or Majorana particles. Furthermore, our framework includes both the neutrino mass and the mediator mass as free parameters. From there, we systematically distill the three approximations used commonly in the literature: the light, heavy, and resonant mediators. In each case, we compared our results with previous results and showed the advantages and limitations of our approach.
As it comes forward from our constructions and findings, the series of initial questions we posed along the introduction to frame the scope of our work do find answers in our present approach. First of all, we have shown along Subsection 4.2 that it is indeed possible to properly compute the collision term of NSI for a wide range of neutrino temperatures using a free mediator mass, at least for the non-resonant interaction regimes, since, in resonant cases, it seems numerically challenging to get reliable results. On the other hand, as already noted, the main differences that Dirac or Majorana-like neutrino nature implies are primarily determined by the available NSI interaction channels (see Figure 1). We have, therefore, adopted a Dirac-like model, since neutrino–neutrino scattering between the same mass state reproduces the HML in the same way as in the Majorana case; nevertheless, one should keep in mind that the neutrino nature should become important when resonances are considered, which lies beyond the scope of this paper. From our results, we would also like to underline that, as expected, neutrino masses can be neglected at sufficiently high temperatures; however, as they become relevant at lower temperatures, they suppress the collision term. Moreover, we should remark that in this work, we have succeeded in providing a careful calculation for thermal corrections on the light mediator case, without using additional approximations, as demonstrated in Sec. 5.
@@
The commonly used HML is a popular framework that has the potential to address cosmological tensions. However, this approach has previously relied on the effective coupling $G_{\rm eff}$ combined with a tight coupling approximation at high temperatures. The latter is particularly well motivated, as previous computations did not permit a systematic exploration of the coupling strength and the mediator mass. In contrast, our work allows exploration of the parameter space using the fundamental coupling $g$ directly, enabling explicit tests of the mediator mass $m_{\varphi}$. We believe this is the most important result of our analysis, and plan to present results with cosmological observables in the future.
@@
Here, we showed that the neutrino mass does not affect the Boltzmann collision term for $m_{\nu}\lesssim 0.06\,{\rm eV}$ for redshifts larger than $1.8\times 10^{4}$. We also relaxed the heavy mediator validity condition down to $m_{\varphi}\geq 10^{2}\,{\rm eV}$, and showed that Dirac-like neutrinos do not exhibit resonance production in $\nu-\nu$ scattering, but they do in the $\nu_{i}-\bar{\nu}_{j}$ scattering. Modeling the resonant regime for Majorana neutrinos remains challenging since it requires the inclusion of one-loop corrections for a complete amplitude description. We anticipate that a full understanding of these corrections will allow for a thorough exploration of heavy mediators in NSI in future work.
In Sec. 5, we presented an exact treatment for the light mediator regime, establishing that for $m_{\varphi}\lesssim 10^{-3}\,{\rm eV}$ the mediator mass and neutrino mass ($m_{j}\lesssim 0.06\,{\rm eV}$) can be safely neglected according to our numerical analysis for $z\gtrsim 6\times 10^{3}$ ($T_{\nu}\gtrsim 1\rm{eV}$). To conclude, we were able to find the correct value of $\xi_{l}$, which is typically codified into $g_{\rm eff}=\xi^{1/4}g$ that was not previously computed, and find the multipolar dependence on $\ell$. All of our results would be useful for future cosmological data analysis of neutrino self-interactions.
@@
Another important feature is that interactions such as (3.1) can produce scalar particles through the process $\varphi$, via $\nu\nu\to\varphi\varphi$ and $\nu\nu\to\varphi$ when the temperature is comparable to $m_{\varphi}$. This may invalidate the assumption $C^{(0)}=0$ (equivalently, $Q^{(0)}=0$), or at least render its contribution subdominant rather than negligible. We plan to address this effect and provide a more accurate description in future work.
@@ Acknowledgments
The metric perturbation $h_{ij}$ can be decomposed into a trace part and a traceless part; The trace part is defined as $h={h^{i}}_{i}$, and the traceless part consists of three components: $h^{(k)}_{ij}$, $h^{(\perp)}_{ij}$, and $h^{(T)}_{ij}$. Thus, we write $h_{ij}=h\,\delta_{ij}/3+h^{(k)}_{ij}+h^{(\perp)}_{ij}+h^{(T)}_{ij}$. By definition, the divergences of $h^{(k)}_{ij}$ and $h^{(\perp)}_{ij}$ correspond to longitudinal and transverse vectors, respectively, while $h^{(T)}_{ij}$ is fully transverse. The components of $h_{ij}$ satisfy the following conditions: $\epsilon_{ijk}\,\partial_{j}\partial^{l}h^{(k)}_{lk}=0$, $\partial^{i}\partial^{j}h^{(\perp)}_{ij}=0\,,\partial^{i}h^{(T)}_{ij}=0\,.$ It follows that $h^{(k)}_{ij}$ can be expressed in terms of a scalar field $\chi$, and $h^{(\perp)}_{ij}$ in terms of a divergenceless vector field $A_{i}$, as
$h^{(k)}_{ij}=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\chi\,,\quad h^{(\perp)}_{ij}=\partial_{i}A_{j}+\partial_{j}A_{i}\,,\quad\partial^{i}A_{i}=0\,.$ (A.1)
@@ Appendix B Collision integral reduction for $\alpha-\beta$ elastic scattering processes
$F^{(1)}=\Psi_{S}\left[2K_{3}\Theta({\bf p}_{3})-K_{1}\Theta({\bf p}_{1})-K_{2}\Theta({\bf p}_{2})\right]\,$, (B.2)
@@
for different particles in the final state. We have dropped the implicit $\tau$ and $\bf x$} dependence; here, the $K_{n}$ terms are determined by
@@
The general form of $\Psi_{S}$ is valid for any elastic scattering processes $\alpha(P_{1})+\beta(P_{2})\rightarrow\alpha(P_{4})+\beta(P_{3})$ and annihilation processes as $\alpha+\alpha\to\beta+\beta$, since the necessary and sufficient condition to establish this expression is $\sigma_{1}\sigma_{2}=\sigma_{3}\sigma_{4}$, though, we will not analyze inelastic processes here  [^15] In certain inelastic scattering processes, the same condition on sigmas applies, such as in $\nu_{i}\nu_{j}\to\varphi\varphi$ or $\nu_{i}\nu_{j}\to\nu_{k}\nu_{l}$.. Note that the double contribution of $\Theta({\bf p}_{3})$ in Eq.(B.2) comes from $\Theta({\bf p}_{4})$ after the explicit use of the ${\bf p}_{4}\leftrightarrow{\bf p}_{3}$ symmetry for identical particles in the final state. The corresponding rescaled collision term in Fourier space then behaves as follows
@@
$\frac{d^{3}p_{4}}{2E_{4}}=d^{4}P_{4}\delta(P_{4}^{2}-m_{j}^{2})H(E_{4}-m_{j})\,,\quad\text{and}\quad\frac{d^{3}p_{3}}{2E_{3}}=d^{4}P_{3}\delta(P_{3}^{2}-m_{i}^{2})H(E_{3}-m_{i})\,.$ (B.11)
@@
$(P_{1}+P_{2}-P_{4})^{2}-m_{i}^{2}$ $=2\left(P_{1}\cdot P_{2}-P_{1}\cdot P_{4}-P_{2}\cdot P_{4}+m_{j}^{2}\right)\,$,
because of Eq. (B.12), fixes the angle to the value $\beta_{0}$ ($\beta_{0}^{\prime}$) that comply with the (respective)condition
@@
$\cos\beta_{0}^{\prime}$ $=\frac{E_{2}E_{4}-E_{1}(E_{2}-E_{4})-m_{j}^{2}+p_{1}p_{2}\cos\alpha-p_{1}p_{4}\cos\theta^{\prime}-p_{2}p_{4}\cos\alpha\cos\theta^{\prime}}{p_{2}p_{4}\sin\alpha\sin\theta^{\prime}}\,$, (B.15)
that we should accordingly for the corresponding case.
@@
$\sin^{2}\beta_{0}=\frac{\tilde{S}(x,y)}{|p_{2}p_{3}\sin\alpha\sin\theta|^{2}}\quad\text{and}\quad\sin^{2}\beta_{0}^{\prime}=\frac{\tilde{S}^{\prime}(x,z)}{|p_{2}p_{4}\sin\alpha\sin\theta^{\prime}|^{2}}\,$, (B.19)
@@
    $c_{x,j}^{\prime}$ $=-[z_{2}z_{4}+z_{1}(z_{4}-z_{2})-\mu_{j}^{2}+q_{1}q_{2}x]^{2}+q_{2}^{2}q_{4}^{2}(1-x^{2})~.$ (B.28)
  • Alternatively, $S^{\prime}(x,z)=a_{z}x^{2}+b_{z,j}x+c_{z,j}$, where
@@
where $I_{n}$ are the angular integrals
$I_{n}=\begin{cases}I_{n}(\mu_{n},p_{1},p_{2},p_{3},\tau)=\frac{1}{T^{2}}\int\frac{dxdy\,d\phi}{\sqrt{S(x,y)}}\Theta_{n}({\bf k},{\bf p}_{n},\tau)\overline{|{\cal M}|}^{2}H[S(x,y)]\quad\text{for $n=1,2,3$}\\ I_{4}(\mu_{4},p_{1},p_{2},p_{4},\tau)=\frac{1}{T^{2}}\int\frac{dxdz\,d\phi^{\prime}}{\sqrt{S^{\prime}(x,z)}}\Theta_{4}({\bf k},{\bf p}_{4},\tau)\overline{|{\cal M}|}^{2}H[S^{\prime}(x,z)]\quad\text{for $n=4$}\end{cases}\,.$ (B.33)
Both expressions were evaluated under the momentum conservation condition $P_{4}=P_{1}+P_{2}-P_{3}$. After inserting the Legendre expansion (2.14), where $k=|{\bf k}|$, and $P_{\ell}$ is the Legendre polynomial of order $\ell$,
@@
${\cal J}_{4}(p_{1},\tau)$ $=\frac{1}{8(2\pi)^{4}E_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1){\cal P}_{\ell}(\upmu)\,\int\!\frac{p_{2}^{2}dp_{2}}{E_{2}}\frac{p_{4}^{2}dp_{4}}{E_{4}}\Psi_{S}K_{4}H(E_{3}-m_{i})\vartheta_{l}(p_{4})\tilde{\cal I}_{4l}\,$,
@@
${\cal I}_{4\ell}(q_{1},q_{2},q_{4})$ $=\int\frac{dxdz}{\sqrt{S^{\prime}}}\overline{|{\cal M}|}^{2}\,H(S^{\prime})\,{\cal P}_{\ell}(z)\,$,
@@
$\tilde{C}^{(1)}[{\bf k},p_{1},\tau]=\frac{1}{4\pi}\frac{T_{\nu}}{128\pi^{3}z_{1}}\frac{d\ln f^{(0)}(p_{1})}{d\ln p_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1)\nu_{\ell}{\cal P}_{\ell}(\upmu)\tilde{{\cal K}}(z_{1})\,$, (B.38)